# Closed Orbits of Dynamically Convex Reeb Flows: Towards the HZ- and Multiplicity Conjectures

### v2

@@ 3.1.2. Dynamical convexity
A word of warning is due: the flow generated by a positive definite Hamiltonian need not be dynamically convex. This is the case for, instance, for the time-one flow of a small Hamiltonian; see [GG20, Ex. 4.7] for further discussion. The proof of the following result is quite standard and can be found in, e.g., [GG20, Sec. 4.2]; cf. [Lo02, SZ].
@@ 3.1.3. Conley–Zehnder index: closed orbits
Let now $M=\partial W$ be the boundary of a Liouville domain $(W^{2n\geq 4},d\alpha)$ such that $c_{1}(TW)=0$ as in Section 2.1 and let $\xi$ be the contact structure $\ker\alpha|_{M}$. Then $c_{1}(\xi)=0$ and hence the top complex wedge power $L_{M}$ of $\xi$ is a trivial complex line bundle over $M$. Likewise, the top complex wedge of $TW$ is a trivial line bundle $L$. Fix a unit section ${\mathfrak{s}}$ of $L$. Note that the symplectic orthogonal to $\xi$ in $T_{M}W$ is trivial and using, say, the Reeb vector field $R$ as a section of this bundle, we have an isomorphism $T_{M}W=\xi\oplus{\mathbb{C}}^{2}$ as bundles over $M$. Let ${\mathfrak{s}}_{M}$ be the unit section of $L_{M}$ such that ${\mathfrak{s}}={\mathfrak{s}}_{M}\otimes R$ over $M$. This section is completely determined by ${\mathfrak{s}}$, up to homotopy.
Next, let $x\colon S_{T}^{1}:={\mathbb{R}}/T{\mathbb{Z}}\to M$ be a closed orbit of the Reeb flow, which we do not require to be prime. The pull-back ${\mathfrak{s}}_{M}\circ x$ of ${\mathfrak{s}}_{M}$ by $x$ is a non-vanishing section of the pull-back $x^{*}L_{M}$ over $S_{T}^{1}$ and $x^{*}L_{M}$ is in turn the top complex wedge of $x^{*}\xi$. Pick any Kähler trivialization of $x^{*}\xi$ which, up to homotopy, gives rise to the trivialization ${\mathfrak{s}}_{M}\circ x$ of $x^{*}L$, i.e., such that its top complex wedge is ${\mathfrak{s}}_{M}\circ x$. (Such a trivialization is unique up to homotopy.) Then we use this trivialization to turn the linearized Reeb flow $D\varphi^{t}$, $t\in[0,T]$, along $x$ into a path $\Phi(t)\in\operatorname{Sp}(2m)$ with $m=n-1$ starting at the identity, and hence an element $\Phi\in\operatorname{\widetilde{Sp}}(2m)$. By definition, $\mu_{\pm}(x):=\mu_{\pm}(\Phi)$ and $\operatorname{\hat{\mu}}(x):=\operatorname{\hat{\mu}}(\Phi)$. In particular, when $x$ is non-degenerate, $\mu(x):=\mu(\Phi)$. Furthermore, $\operatorname{\hat{\mu}}$ is homogenenous, i.e.,
@@ 3.2. Index recurrence and applications
The original version of the index recurrence theorem from [GG20] asserts that arbitrarily long intervals in the sequence of the Conley–Zehnder indices of the iterates of one or a finite collection of symplectic linear maps repeat themselves infinitely many times up to a shift. A more general variant of the theorem which we need here allows for control of not only the indices but also additional real parameters assigned to the maps, which should be thought of as periods of closed orbits. In addition, it shows that the sequence of “recurrence events” can be made to have a bounded gap.
@@
In what follows we will to refer to the collection $\big{\{}k_{ijl},d_{il},C_{l}\big{\}}$ for a fixed $l$ as a recurrence event.
@@
$\begin{split}\mu_{-}(\Phi^{k}_{ij})-d_{il}&\geq\big{(}\operatorname{\hat{\mu}}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})\big{)}-\big{|}\mu_{-}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k}_{ijl})\big{|}-\big{|}\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})-d_{il}\big{|}\\ &\geq(k-k_{ijl})\operatorname{\hat{\mu}}(\Phi_{ij})-m-\eta\\ &\geq 2(k-k_{ijl})-m-1\\ &\geq 2m+2.\end{split}$
@@
$\begin{split}\mu_{+}(\Phi^{k}_{ij})-d_{l}&\leq\big{(}\operatorname{\hat{\mu}}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})\big{)}+\big{|}\mu_{-}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k}_{ijl})\big{|}+\big{|}\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})-d_{il}\big{|}\\ &\leq(k-k_{ijl})\operatorname{\hat{\mu}}(\Phi_{ij})+m+\eta\\ &\leq 2(k-k_{ijl})+m+1\\ &\leq-2m-2.\end{split}$
@@
$\begin{split}\mu_{+}\big{(}\Phi^{k}_{ij}\big{)}&=d_{il}-\mu_{+}\big{(}\Phi^{\ell}_{ij}\big{)}+\big{(}\beta_{+}\big{(}\Phi^{\ell}_{ij}\big{)}-\beta_{+}\big{(}\Phi^{\ell}_{ij}\big{)}\big{)}\\ &\leq d_{il}-m-2+m\\ &=d_{il}-2\end{split}$
by (IR3) and again the dynamical convexity assumption, proving (IR2’). When $\Phi_{ij}$ is strongly non-degenerate, the term $\beta_{+}-\beta_{-}$ disappears and we obtain the upper bound $d_{il}-m-2$. ∎
@@
In the setting of Theorem 3.3 or Corollary 3.5, for a collection of maps $\Phi_{j}\in\widetilde{\operatorname{Sp}}(2m)$, we can find bounded gap sequences $C_{l}$ and $d_{l}$ independent of $j$, and bounded gap sequences $k_{jl}$ such that all other conditions are satisfied. In particular, the sequences $k_{jl}$ and $d_{l}$ can be made divisible by any fixed integer. Furthermore, with $\Delta_{j}:=\operatorname{\hat{\mu}}(\Phi_{j})$ condition (IR5) takes the following form: $C_{l}-\eta<k_{jl}\Delta_{j}<C_{l}$, and $k\Delta_{j}<C_{l}-\eta$ when $k<k_{ijl}$ and $k\Delta_{ij}>C_{l}$ when $k>k_{jl}$.
@@
This corollary is close to the original version of the index recurrence theorem from [GG20, Thm. 5.2], which in turn is related to and can be derived from the common jump theorem from [Lo02, LZ]. The new point of Corollary 3.6 as compared to these results is the part that all sequences involved have bounded gap. A simple proof from [GG20] of the following multiplicity result, quite standard now, is an illustrative application of the index recurrence theorem in the form of Corollaries 3.5 and 3.6. Interestingly, the argument relies on the existence of just one recurrence event.
@@
Without loss of generality, we may assume that there are only finitely many prime closed orbits $x_{i}$ in $M$. Denote the linearized flows along these orbits by $\Phi_{1},\ldots,\Phi_{r}$. Consider one recurrence event from Corollary 3.6 for these elements of $\operatorname{Sp}(2(n-1))$. Suppressing $l$ in the notation and ignoring $C_{l}$, we denote this event by $\{d,k_{1},\ldots,k_{r}\}$ and require $d$ and all $k_{j}$ to be divisible by two. Since $\operatorname{CH}_{n-1+2i}(W;{\mathbb{Q}})\neq 0$, $i\in{\mathbb{N}}$, there must be at least one closed orbit with index $n-1+2i$. There are exactly $n$ such integer points in the range ${\mathcal{I}}:=[d-n+1,d-n-1]$. By Corollary 3.5, the only closed Reeb orbits with index in ${\mathcal{I}}$ are $x_{j}^{k_{j}}$. Moreover, since $k_{j}$ is even, to contribute to the symplectic homology, $x_{j}$ must be non-alternating and hence, the parity of $\mu(x_{j})$ must be the same as that of $\mu\big{(}x_{j}^{k_{j}}\big{)}$, i.e., $n+1$. Therefore, there are at least $n$ non-alternating prime closed orbits with parity $n+1$. ∎
@@
As we have already mentioned in the introduction, Corollary 3.7 had been generalized and strengthened in several directions. When the non-degeneracy condition is removed, a proof similar to the above one but relying on equivariant spectral invariants gives a lower bound $\lfloor n/2\rfloor+1$. The argument can be further pushed (in a non-trivial way) to improve this lower bound to $\lceil n/2\rceil+1$, [DL, GG20]. In dimensions 6 and 8, the lower bound is 3; see [Wa16a, WHL].) Furthermore, for non-degenerate flows, the dynamical convexity condition has been replaced by a much less restrictive index requirement, [DL2W18, DL2W22, GGM]. Corollary 3.8 has also been refined and generalized in a variety of ways; see, e.g., [AM22, ALM, GiM] and references therein.
@@ 4. Proofs of the main results
Here we first derive Theorems A and B from Theorem E in Section 4.1. Then, in Section 4.2, we establish the latter as a consequence of Theorem C. In Section 4.3, we turn to the proofs of Theorem C, which is the central part of the argument, and Corollary D. Finally, Corollary F is proved in Section 4.4
@@ 4.1. Proofs of Theorems A and B
4.1.1. Proof of Part (M1) of Theorem A
It is convenient to break down the proof into three steps.

### v3

@@ 3.1.2. Dynamical convexity
A word of warning is due: the flow generated by a positive definite Hamiltonian need not be dynamically convex. This is the case for, instance, for the time-one flow of a small Hamiltonian; see [GG20, Ex. 4.7] for further discussion. The proof of the following result is quite standard and can be found in, e.g., [Lo02, SZ] and also [GG20, Sec. 4.2].
@@ 3.1.3. Conley–Zehnder index: closed orbits
Now let $M=\partial W$ be the boundary of a Liouville domain $(W^{2n\geq 4},d\alpha)$ such that $c_{1}(TW)=0$ as in Section 2.1 and let $\xi$ be the contact structure $\ker\alpha|_{M}$. Then $c_{1}(\xi)=0$ and hence the top complex wedge power $L_{M}$ of $\xi$ is a trivial complex line bundle over $M$. Likewise, the top complex wedge of $TW$ is a trivial line bundle $L$. Fix a unit section $\mathfrak{s}$} of $L$. Note that the symplectic orthogonal to $\xi$ in $T_{M}W$ is trivial and using, say, the Reeb vector field $R$ as a section of this bundle, we have an isomorphism $T_{M}W=\xi\oplus{\mathbb{C}}$ as bundles over $M$. Let ${\mathfrak{s}}_{M}$ be the unit section of $L_{M}$ such that ${\mathfrak{s}}={\mathfrak{s}}_{M}\otimes R$ over $M$. This section is completely determined by $\mathfrak{s}$}, up to homotopy.
Next, let $x\colon S_{T}^{1}:={\mathbb{R}}/T{\mathbb{Z}}\to M$ be a closed orbit of the Reeb flow, which we do not require to be prime. The pull-back ${\mathfrak{s}}_{M}\circ x$ of ${\mathfrak{s}}_{M}$ by $x$ is a non-vanishing section of the pull-back $x^{*}L_{M}$ over $S_{T}^{1}$ and $x^{*}L_{M}$ is in turn the top complex wedge of $x^{*}\xi$. Pick any Kähler trivialization of $x^{*}\xi$ which, up to homotopy, gives rise to the trivialization ${\mathfrak{s}}_{M}\circ x$ of $x^{*}L$, i.e., such that its top complex wedge is ${\mathfrak{s}}_{M}\circ x$. (Such a trivialization is unique up to homotopy.) Then we use this trivialization to turn the linearized Reeb flow $D\varphi^{t}$, $t\in[0,T]$, along $x$ into a path $\Phi(t)\in\operatorname{Sp}(2m)$ with m=n1m=n-1 starting at the identity, and hence an element $\Phi\in\operatorname{\widetilde{Sp}}(2m)$. By definition, $\mu_{\pm}(x):=\mu_{\pm}(\Phi)$ and $\operatorname{\hat{\mu}}(x):=\operatorname{\hat{\mu}}(\Phi)$. In particular, when $x$ is non-degenerate, $\mu(x):=\mu(\Phi)$. Furthermore, $\operatorname{\hat{\mu}}$ is homogeneous, i.e.,
@@ 3.2. Index recurrence and applications
The original version of the index recurrence theorem from [GG20] asserts that arbitrarily long intervals in the sequence of the Conley–Zehnder indices of the iterates of one or a finite collection of symplectic linear maps repeat themselves infinitely many times up to a shift. (That theorem can also be derived in a non-trivial way from the index iteration formula, aka the common jump theorem, in [LZ]; we will return to the relation between the two results later in this section.) A more general variant of the theorem which we need here allows for control of not only the indices but also additional real parameters assigned to the maps, which should be thought of as periods of closed orbits. In addition, it shows that the sequence of “recurrence events” can be made to have a bounded gap.
@@
The statement of our next result is admittedly notation heavy. It might be helpful for the reader to keep in mind that in what follows $l$ indexes recurrence events, $k$ with various subscripts and $\ell$ stand for the order of iteration, and $d_{il}$ is the index shift, i.e., an integer approximating the mean indices of the iterates $\Phi^{k_{ijl}}_{ij}$.
@@
In what follows we will refer to the collection $\big\{k_{ijl},d_{il},C_{l}\big\}$ for a fixed $l$ as a recurrence event.
@@
$\begin{split}\mu_{-}(\Phi^{k}_{ij})-d_{il}&\geq\big(\operatorname{\hat{\mu}}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})\big)-\big|\mu_{-}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k}_{ij})\big|-\big|\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})-d_{il}\big|\\ &\geq(k-k_{ijl})\operatorname{\hat{\mu}}(\Phi_{ij})-m-\eta\\ &\geq 2(k-k_{ijl})-m-1\\ &\geq 2m+2.\end{split}$
@@
$\begin{split}\mu_{+}(\Phi^{k}_{ij})-d_{il}&\leq\big(\operatorname{\hat{\mu}}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})\big)+\big|\mu_{+}(\Phi^{k}_{ij})-\operatorname{\hat{\mu}}(\Phi^{k}_{ij})\big|+\big|\operatorname{\hat{\mu}}(\Phi^{k_{ijl}}_{ij})-d_{il}\big|\\ &\leq(k-k_{ijl})\operatorname{\hat{\mu}}(\Phi_{ij})+m+\eta\\ &\leq 2(k-k_{ijl})+m+1\\ &\leq-2m-2.\end{split}$
@@
$\begin{split}\mu_{+}\big(\Phi^{k}_{ij}\big)&=d_{il}-\mu_{-}\big(\Phi^{\ell}_{ij}\big)+\big(\beta_{+}\big(\Phi^{\ell}_{ij}\big)-\beta_{-}\big(\Phi^{\ell}_{ij}\big)\big)\\ &\leq d_{il}-m-2+m\\ &=d_{il}-2\end{split}$
by (IR3) and again the dynamical convexity assumption, proving (IR3’). When $\Phi_{ij}$ is strongly non-degenerate, the term $\beta_{+}-\beta_{-}$ disappears and we obtain the upper bound $d_{il}-m-2$. ∎
@@
In the setting of Theorem 3.3 or Corollary 3.5, for a collection of maps $\Phi_{j}\in\widetilde{\operatorname{Sp}}(2m)$, we can find bounded gap sequences $C_{l}$ and $d_{l}$ independent of $j$, and bounded gap sequences $k_{jl}$ such that all other conditions are satisfied. In particular, the sequences $k_{jl}$ and $d_{l}$ can be made divisible by any fixed integer. Furthermore, with $\Delta_{j}:=\operatorname{\hat{\mu}}(\Phi_{j})$ condition (IR5) takes the following form: $C_{l}-\eta<k_{jl}\Delta_{j}<C_{l}$, and $k\Delta_{j}<C_{l}-\eta$ when $k<k_{jl}$ and $k\Delta_{j}>C_{l}$ when $k>k_{jl}$.
@@
This corollary is close to the original version of the index recurrence theorem from [GG20, Thm. 5.2]. This theorem is in turn closely related to and can be derived from the common jump theorem from [Lo02, LZ] or from its further refined variant proved in [DLW], which is done in detail in [DLWZ]. The new point of Corollary 3.6 as compared to these results is the part that all sequences involved have bounded gap. A simple proof from [GG20] of the following multiplicity result, quite standard now, is an illustrative application of the index recurrence theorem in the form of Corollaries 3.5 and 3.6. Interestingly, the argument relies on the existence of just one recurrence event.
@@
Without loss of generality, we may assume that there are only finitely many prime closed orbits $x_{i}$ in $M$. Denote the linearized flows along these orbits by $\Phi_{1},\ldots,\Phi_{r}$. Consider one recurrence event from Corollary 3.6 for these elements of $\operatorname{Sp}(2(n-1))$. Suppressing $l$ in the notation and ignoring $C_{l}$, we denote this event by $\{d,k_{1},\ldots,k_{r}\}$ and require $d$ and all $k_{j}$ to be divisible by two. Since $\operatorname{CH}_{n-1+2i}(W;{\mathbb{Q}})\neq 0$, $i\in{\mathbb{N}}$, there must be at least one closed orbit with index n1+2in-1+2i. There are exactly $n$ such integer points in the range ${\mathcal{I}}:=[d-n+1,d+n-1]$. By Corollary 3.5, the only closed Reeb orbits with index in $\mathcal{I}$} are $x_{j}^{k_{j}}$. Moreover, since $k_{j}$ is even, to contribute to the symplectic homology, $x_{j}$ must be non-alternating and hence, the parity of $\mu(x_{j})$ must be the same as that of $\mu\big(x_{j}^{k_{j}}\big)$, i.e., $n+1$. Therefore, there are at least $n$ non-alternating prime closed orbits with parity $n+1$. ∎
@@
As we have already mentioned in the introduction, Corollary 3.7 had been generalized and strengthened in several directions. When the non-degeneracy condition is removed, a proof similar to the above one but relying on equivariant spectral invariants gives a lower bound $\lfloor n/2\rfloor+1$. The argument can be further pushed (in a non-trivial way) to improve this lower bound to $\lceil n/2\rceil+1$; see [Wa16b] and also [DL, GG20]. In dimensions 6 and 8, the lower bound is 3 and, respectively 4; see [DL, Wa16a, WHL].) Furthermore, for non-degenerate flows, the dynamical convexity condition has been replaced by a much less restrictive index requirement, [DL2W18, DL2W24, GGM]. Corollary 3.8 has also been refined and generalized in a variety of ways; see, e.g., [AM22, ALM, GiM] and references therein.
@@ 4. Proofs of the main results
Here we first derive Theorems A and B from Corollary D and Theorem E in Section 4.1. Then, in Section 4.2, we establish the latter as a consequence of Theorem C. In Section 4.3, we turn to the proofs of Theorem C, which is the central part of the argument, and Corollary D. Finally, Corollary F is proved in Section 4.4
@@ 4.1. Proofs of Theorems A and B
4.1.1. Proof of Theorem A
Our proof of the theorem actually establishes a more precise result asserting the existence of exactly $n$ prime orbits satisfying certain visibility conditions. Although this result is more technical, we state and prove it here, for it sheds more light on the proof and is of independent interest.
Let $y$ be a prime orbit and let $p(y)$ be the least common multiple of the degrees of the roots of unity among the elliptic eigenvalues of the return map of $y$. Thus $y^{p(y)}$ is a “maximally degenerate” iterate of $y$. Note that $y^{p(y)}$ can fail to be $\mathbb{Q}$}-invisible even when $y$ is elliptic and non-degenerate; see Example 2.12. Set $N(y):=2p(y)$. We will call $y$ virtually $\mathbb{Q}$}-visible when $y^{N(y)}$ is $\mathbb{Q}$}-visible. For instance, a strongly non-degenerate prime orbit is virtually $\mathbb{Q}$}-visible if and only if it is non-alternating. A totally degenerate $\mathbb{Q}$}-visible prime orbit is virtually $\mathbb{Q}$}-visible. The orbit from Example 2.12 is not. As of writing, it is not known if a virtually $\mathbb{Q}$}-visible prime orbit must be $\mathbb{Q}$}-visible even in the setting of Theorem A.
Theorem 4.1.
Assume that the Reeb flow on the boundary $M^{2n-1}\subset{\mathbb{R}}^{2n}$ of a star-shaped domain is dynamically convex and has finitely many prime closed orbits. Then the flow has exactly $n$ virtually $\mathbb{Q}$}-visible prime orbits. Moreover, these orbits $y$ have the same ratio ${\mathcal{A}}(y)/\operatorname{\hat{\mu}}(y)$.
Remark 4.2.
The $n$ virtually $\mathbb{Q}$}-visible closed orbits from Theorem 4.1 have an additional property that $\operatorname{\hat{\mu}}(y)/\operatorname{\hat{\mu}}(y^{\prime})\not\in{\mathbb{Q}}$ for any two such distinct orbits $y$ and $y^{\prime}$. Indeed, assume that this ratio is rational: $\operatorname{\hat{\mu}}(y)/\operatorname{\hat{\mu}}(y^{\prime})=p/q$ for some positive integers $p$ and $q$. Let $m$ be a common multiple of $N(y)$ and $N(y^{\prime})$. Then, by (4.1) below, $y^{qm}$ and $(y^{\prime})^{pm}$ are both $\mathbb{Q}$}-visible. The iterates $y^{qm}$ and $(y^{\prime})^{pm}$ have the same action (aka period), which contradicts Corollary D. This observation, added after we have learned about [Wa25], sharpens the main result from that paper, which in turn is a refinement of a theorem from [HO].
Part (M1) of Theorem A immediately follows from Theorem 4.1. To prove (M2), recall that in the non-degenerate case a prime orbit is virtually $\mathbb{Q}$}-visible if and only if it is non-alternating. Hence there are exactly $n$ such orbits. The parity statement is a consequence Part (O1) from Theorem E since now the degree is equal to the Conley–Zehnder index. As we have already mentioned, the lower bound in (M2) is not new; cf. Corollary 3.7 and [GG20, GK, LZ].
Proof of Theorem 4.1.
The “Moreover” part follows immediately from (O2) in Theorem E. We begin with some general remarks on virtually $\mathbb{Q}$}-visible orbits. Then we turn to the multiplicity statement which comprises the main part and the bulk of the proof.
Let $y$ be a closed orbit. Then, in the setting of the theorem,
[…]
In other words, $y$ is virtually $\mathbb{Q}$}-visible if and only if there exists $k\in{\mathbb{N}}$ such that $y^{kN(y)}$ is $\mathbb{Q}$}-visible, and if and only if $y^{kN(y)}$ is $\mathbb{Q}$}-visible for all $k\in{\mathbb{N}}$. (In particular, again within the theorem’s framework, for any orbit $z$ which is not virtually $\mathbb{Q}$}-visible, no iterate $z^{kN(y)}$, $k\in{\mathbb{N}}$, is $\mathbb{Q}$}-visible.)
Indeed, $y^{kN(y)}$ is an admissible iterate of $y^{N(y)}$ and the return map of $y^{N(y)}$ has no real negative eigenvalues by the definition of $N(y)$. Hence,
$\chi^{\scriptscriptstyle{eq}}\big(y^{kN(y)}\big)=\chi^{\scriptscriptstyle{eq}}\big(y^{N(y)}\big)$
for all $k\in{\mathbb{N}}$ by Lemma 2.8. Under the conditions of the theorem, $\chi^{\scriptscriptstyle{eq}}(x)$ is either $(-1)^{n+1}$ or 0 depending on whether $x$ is $\mathbb{Q}$}-visible or not, as follows from Corollary D. Hence, $y^{N(y)}$ and $y^{kN(y)}$ are both either $\mathbb{Q}$}-visible or $\mathbb{Q}$}-invisible.
The strategy of the proof of the multiplicity part is to use the index recurrence theorem to identify an interval ${\mathcal{I}}:=[d-n+1,d+n-1]\subset{\mathbb{N}}$ such that for each eventually $\mathbb{Q}$}-visible prime orbit $x_{j}$ only one iterate $x_{j}^{k_{j}}$ can have degree in $\mathcal{I}$}. Moreover, we ensure that $x_{j}^{k_{j}}$ is $\mathbb{Q}$}-visible if and only if $x_{j}$ is virtually $\mathbb{Q}$}-visible and thus there are exactly $n$ such orbits. The main difficulty lies in showing that $\deg(x_{j}^{k<k_{j}})\not\in{\mathcal{I}}$ and this is where we crucially rely on Corollary D and Theorem E. We prove this first for the least action orbit $x$ by analyzing the behavior of the sequence $\deg(x^{k})$. Dealing with the iterates of the remaining prime orbits $x_{j}$, we show that $x_{j}^{k}$ must be $\mathbb{Q}$}-invisible whenever $\deg(x_{j}^{k})\in{\mathcal{I}}$ and $k<k_{j}$, by comparing the degree and the action of $x_{j}^{k}$ with the degree and the action of the iterates of $x$. It is convenient to break down the proof into three steps.